8. A 2.71 g sample of KCl is found to be radioactive. It is decaying at a rate of 4490


Bq. The decay is traced to an isotope of potassium 40K, which constitutes 1.17% of


normal potassium. Calculate the half life of this nuclide.



1
Expert's answer
2021-12-27T08:19:00-0500

"A=\eta \frac m M N_A\frac{\ln 2}{T^{1/2}},\implies"

"T^{1/2}=\frac{\eta mN_A \ln2}{AM}=1.25\cdot 10^9~"years.